1. A charged parallel plate with 15F connected to a battery with 9V. When a dielectric material was inserted between the parallel plates. What happens to the stored charges in the capacitor? 2. Find the capacitance of a layer of Al203 that is 0.5 m thick and 2000 mm2 of square area ( r= 8.854 x 10-12FM-1).

1
Expert's answer
2022-04-11T12:44:29-0400

1. The charge on the capacitors plates increases.

2.

"C=\frac{\kappa\epsilon_0 A}{d}"

"C=\frac{9*8.85*10^{-12}*2000*10^{-6}}{0.5}=3.2*10^{-13}\:\rm F"